OpenStudy (anonymous):

Find the exact value of sin75 using any appropriate formulas. help please :D

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

sin75 is the same as sin(45+30)

OpenStudy (anonymous):

from there you do sin30*cos45+cos30*sin45

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

which if you find those coordinates = \[1/2 * \sqrt{2}/2 + \sqrt{3}/2 * \sqrt{2}/2\]

OpenStudy (anonymous):

= \[\sqrt{2}/4+\sqrt{6}/4\]

OpenStudy (anonymous):

ty so much

OpenStudy (anonymous):

can u help with another question?

OpenStudy (anonymous):

combine you get \[(\sqrt{2} + \sqrt{6})/\]

OpenStudy (anonymous):

/4

OpenStudy (anonymous):

sure thing and ya i can help

OpenStudy (anonymous):

find the exact value of [\sec ^{2}(2\pi/3)+\tan(5\pi/4)(\cos \pi/3)\]

OpenStudy (anonymous):

ah crap let me re write it

OpenStudy (anonymous):

haha sure thing

OpenStudy (anonymous):

\[\sec ^{2}(2\pi/3)+\tan(5\pi/4)(\cos \pi/3)\]

OpenStudy (anonymous):

finally lol

OpenStudy (anonymous):

haha ok gimme a sec to figure out how to do it and then how to explain it XD

OpenStudy (anonymous):

no problem :D

OpenStudy (anonymous):

are you trying to find the # value or just simplify it?

OpenStudy (anonymous):

because sec^2(2pi/3) = 4

OpenStudy (anonymous):

it says find the exact value

OpenStudy (anonymous):

i know the answer but idk how to get it

OpenStudy (anonymous):

the answer is 4.5

OpenStudy (anonymous):

ok well then the other half =.5 then XD ok.....so to solve the first part....gimme sec to type this out

OpenStudy (anonymous):

sec^2(2pi/3) is the same as saying (sec(2pi/3))^2

OpenStudy (anonymous):

would i have to jsut type all this into a calculator?

OpenStudy (anonymous):

that you can then change to (1/cos(pi/3)^2

OpenStudy (anonymous):

im sure you could but its fine, i can just explain it here

OpenStudy (anonymous):

cos(pi/3)= 1/2 so you have (1/1/2)^2 = (2)^2 = 4

OpenStudy (anonymous):

yea i think i understand this now haha

OpenStudy (anonymous):

i got one more if you are up for it :D

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

solve for x to the nearest degree over 0<x<360 in the equation \[4\cos x-3secx=2\]

OpenStudy (anonymous):

you are suppose to multiply by cosx right?

OpenStudy (anonymous):

and get \[4\cos ^{2}x-3=\]

OpenStudy (anonymous):

2cosx

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i get stuck here on what to do next

OpenStudy (anonymous):

ok...let me see.....

OpenStudy (anonymous):

here, one sec ill be right back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so, im double checking, but i got undefined as my answer....

OpenStudy (anonymous):

the answer key says its 131 and 229

OpenStudy (anonymous):

but i coudlnt get an answer either so idk

OpenStudy (anonymous):

sorry.